# PAFUSE: Part-based Diffusion for 3D Whole-Body Pose Estimation

### v1

@@ 3 Method
Conditioned inference. Test time inference predicts $\hat{\mathrm{y}}$ by concatening the 2D input tensor $\mathrm{x}$ to a pure-noise tensor sampled from a Gaussian distribution $\hat{\mathrm{y}}^{T}\sim\mathcal{N}(0,I)$. Although the training objective is one-step denoising (Eq. 1), inference still requires incremental denoising ($\hat{\mathrm{y}}^{T}\rightarrow\hat{\mathrm{y}}^{T-1}\rightarrow\cdots\rightarrow\hat{\mathrm{y}}$) to produce high-quality samples from the expected conditional distribution. Following D3DP, we employ DDIM [46] to bridge the gap between the one-step results produced by the learned denoiser and the high-quality incremental denoising needed as final results.
@@ 4.3 Ablations
The part-based loss of Eq. 2 brings the most dramatic gains, in contrast to the whole-body loss of Eq. 3, demonstrating that separating the optimization targets for each part results in substantial improvement. Balancing the size of the denoiser models (384, 256 and 224 channels for body, hands, and face respectively) also considerably improved the results in comparison to using an uniform-sized (288 channels) denoiser.
Other design choices brought smaller, but still visible improvements. N=27 $\rightarrow$ N=81 experiment shows that extending the temporal context adversely affects performance. We attribute this decline to the uneven sampling characteristics of the H3WB dataset. The ablation study comparing MPJPE to MSE shows that despite the common use of MSE loss in diffusion-based methods for 3D whole-body pose estimation, optimizing with MPJPE yields superior results. This outcome aligns with expectations, given that model performance is evaluated using the MPJPE metric.
@@ 5 Conclusion
Appendix 0.A Appendix
We supplement the main article with details on how we adaptated the spatio-temporal models (VideoPose [37], PoseFormer [63], MixSTE [61] and D3DP [43]) to work with our part-based approach.
Technical details on part-based adaption of spatio-temporal models.
In order to support VideoPose [37], PoseFormer [63], MixSTE [61] and D3DP [43] for part-based 3D whole-body pose estimation, we adapt each model as follows:
  1. 1.
    We create 3 pose estimation networks for body, face and hands, and adapt each network to accept the number of keypoints in their corresponding parts.
  2. 2.
    Next, we train each network with the part-shifted keypoints on the corresponding body parts.
  3. 3.
    In order to preserve the original number of parameter for the models, we adjust the number of the input channels in the first layer of each network for each body model. For instance, in VideoPose [37], the number of channels of the first layer is $1024$, yielding $9.6M$ trainable parameters for whole-body pose estimation (directly with 133 keypoints). In the part-based model, we set number of channels to $1024$, $256$, $256$ for the body, face, and hand networks, respectively, yielding $9.9M$ trainable parameters. We present detailed comparisons for each model in Table 4.
Model WB Embed. Dim. Size PB Embed. Dim. Size body / face / hand WB Num. Params PB Num. Params
[…]
Table 4: Detailed comparison of whole-body and part-based pose estimation models, showing input layer embedding dimension and total number of parameters. WB: Whole-Body. PB: Part-Based.

### v2

@@ 3 Method
Conditioned inference.
Test time inference predicts $\hat{\mathrm{y}}$ by concatening the 2D input tensor $\mathrm{x}$ to a pure-noise tensor sampled from a Gaussian distribution $\hat{\mathrm{y}}^{T}\sim\mathcal{N}(0,I)$. Although the training objective is one-step denoising (Eq. 1), inference still requires incremental denoising ($\hat{\mathrm{y}}^{T}\rightarrow\hat{\mathrm{y}}^{T-1}\rightarrow\cdots\rightarrow\hat{\mathrm{y}}$) to produce high-quality samples from the expected conditional distribution. Following D3DP, we employ DDIM [46] to bridge the gap between the one-step results produced by the learned denoiser and the high-quality incremental denoising needed as final results.
@@ 4.3 Ablations
The part-based loss of Eq. 2 brings the most dramatic gains, in contrast to the whole-body loss of Eq. 3, demonstrating that separating the optimization targets for each part results in substantial improvement. Balancing the size of the denoiser models (384, 256 and 224 channels for body, hands, and face respectively) also considerably improved the results in comparison to using an uniform-sized (288 channels) denoiser. Other design choices brought smaller, but still visible improvements. N=27 $\rightarrow$ N=81 experiment shows that extending the temporal context adversely affects performance. We attribute this decline to the uneven sampling characteristics of the H3WB dataset. The ablation study comparing MPJPE to MSE shows that despite the common use of MSE loss in diffusion-based methods for 3D whole-body pose estimation, optimizing with MPJPE yields superior results. This outcome aligns with expectations, given that model performance is evaluated using the MPJPE metric.